OpenStudy (anonymous):

Logistic parent function: 1/1+e^-x. What is the graph and function if 1/+e^-x is shifted left 1 unit? What is the graph and function if flipped over the x-axis?

OpenStudy (anonymous):

\[f(x) = \frac{ 1 }{ 1+e^{-x} }\]

OpenStudy (solomonzelman):

when you shift a function C units right, you subtract C from every x. when you shift C units left, you add C to every x. --------------------------------------------- To reflect the function across the x-axis, mutliply the [entire] function times 1.

OpenStudy (anonymous):

So would the function for shifting left one unit horizontally be: \[f(x) = 1/(1+e^x) + 1\]

OpenStudy (solomonzelman):

that would be shifted 1 unit up.

OpenStudy (anonymous):

The one being under the fraction as well.

OpenStudy (anonymous):

Sorry!

OpenStudy (solomonzelman):

you add/subtract the C not from/to the function. you are subtracting/adding to/from the x.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=\frac{1 }{1+e^{-x}} }\) to shift to the left by 1 unit: \(\large\color{black}{ \displaystyle f(x)=\frac{1 }{1+e^{-(x+1)}} }\)

OpenStudy (solomonzelman):

you are adding C to the x (jut as I did). And in this case, the "C" is 1.

OpenStudy (anonymous):

I see! How would the over x axis function look?

OpenStudy (solomonzelman):

you just multiply the entire function times -1, to reflcet over the x-axis.

OpenStudy (solomonzelman):

to reflect*

OpenStudy (anonymous):

So: \[f(x) = -1(\frac{ 1 }{ 1+e^-x }\]

OpenStudy (anonymous):

Forgot the end parenthesis.

OpenStudy (solomonzelman):

Yes, and that would be just same as \(\displaystyle \large \frac{-1}{1+e^{-\left(x+1\right)}}\)

OpenStudy (anonymous):

Wow! Thank you, I understand now.

OpenStudy (solomonzelman):

Anytime!